# Shoal++: High Throughput DAG BFT Can Be Fast!

### v1

@@ 4. Shoal++ Overview
More anchors!
Next, Shoal++ takes Shoal’s idea of increased anchor frequency one step further, and tries to turn as many nodes into anchors as possible. Doing so requires striking balance. On the one hand, increasing the number of anchors increases the opportunities for anchoring latency that we can avoid. On the other hand, it introduces also opportunity for more uncommitted anchors, which can stall progress. Adopting the reputation scheme proposed by Shoal can help identify a reliable subset of candidates, but is not enough to streamline progress in practice. Shoal++ (i) slows down round advancement to put the DAG into lockstep and avoid missed anchors, and (ii) dynamically re-interprets the tentative anchor schedule to skip anchors that are no longer needed.
More DAGs!
Finally, Shoal++ minimized Queuing latency by operating not one, but multiple DAGs in parallel. DAGs are staggered using a small offset, allowing waiting transactions that miss a round to quickly be included in the next DAG. This simultaneously improves throughput via better resource utilization: proposals are sent more frequently, batches are smaller and can be sent and processed more efficiently (in a streaming fashion) than less frequent, large batches.
5. Shoal++ In-Depth
We now describe the full Shoal++ system in detail.
Shoal++ uses as starting points the DAG-core of Narwahl (Danezis et al., 2022), and Consensus-core of Bullshark (Spiegelman et al., 2022a) outlined in §3. For the sake of brevity, we do not duplicate the pseudocode that is available in previous works and, in particular, omit a detailed description of the DAG construction (i.e., node broadcasts, transaction submission, node fetching, garbage collection, and so on). We defer to the original papers for an in-depth discussion and correctness proofs.
Algorithm 1 summarizes the DAG API and node structure. We abstract away the node certification process, and the commit logic. Process_node and Process_certified_node respectively involve checking that a node proposal is valid and voting for it, and adding a new confirmed node to the DAG, prompting invocation of commit rules. Run_bullshark performs an instance of the Bullshark commit logic to determine (given a starting point) the first anchor to be ordered (§3.1.1).
Upon committing an anchor, all of its (yet-to-be-ordered) causal history is deterministically ordered as well. Helper functions causal_history and get_anchors aid, respectively, in identifying the causal history of a node (necessary for both commit and ordering logic), and the anchor candidate schedule. Shoal++ adopts and extends Shoal’s (Spiegelman et al., 2023) reputation logic; get_anchors returns, for each round, a vector of eligible anchors.
@@ 5. Shoal++ In-Depth
4:     causal_history(v: Node) $\rightarrow$ Vec¡Node¿
@@
6:     get_anchors(round: int) -¿ Vec¡Node¿
@@
3:     weak_votes: Vec¡Vec¡int¿¿ $\triangleright$ Initially empty
@@
5:     anchors: Vec¡Node¿ $\triangleright$ Initially empty
@@
5.1. A faster Direct Commit rule
@@
Figure 3. Bullsharks’s direct commit rule requires f+1 certified vote nodes, while Shoal++’s direct commit rule requires $12f+$ 1 (uncertified) node proposals.
@@
We note, however, that this rule places a stronger requirement for liveness; $f$ additional proposals are (tentatively) required to link to an anchor. We thus retain the existing Direct Commit rule as backup, and allow replicas to commit using whichever rule is satisfied first. In some cases it may be faster (and more reliable) to commit using $f+1$ certified nodes. For instance when the network is unstable, or point latencies are asymmetric, the fastest $f+1$ replicas might advance much more quickly than the fastest $12f+$ 1.
@@
In our empirical evaluation we find that this case is rare (§8) and that replicas are, most of the time, able take advantage of the Fast Direct Commit rule. Shoal’s reputation scheme ensures that eligible anchor candidates are typically well connected, making it exceedingly commonplace for future round proposals to link to an anchor.
5.2. Increasing anchor frequency
@@
In order to reduce Anchoring latency, Shoal++ tries to dynamically designate as many nodes as possible as anchors. Intuitively, if all nodes were anchors that become committed, no node would experience any Anchoring latency. [^3] We note that such an approach is briefly discussed in the closing remarks of Shoal (Spiegelman et al., 2023), but is neither implemented nor carefully evaluated for practical considerations. Safely operating more than one anchor per round requires additional orchestration. To ensure a consistent log order across all correct replicas, parallel anchors must be committed in a pre-defined order. One easy way to visualize parallel orchestration is by mapping anchors to fixed slots; much like parallel proposals in traditional BFT protocols such as PBFT (Castro and Liskov, 2002). Figure 4 shows a simple example in which there are four anchors in one round (all $n$ nodes are anchors). Each anchor employs it’s own consensus instance that may proceed in parallel, and out of order, but anchors can be committed only in sequence. Each consensus instance operates a one-shot Bullshark instance, augmented with the Fast Direct Commit Rule. We remark that each Bullshark instance operates it’s own tentative anchor schedule (e.g. the yellow anchor in Fig. 4 is considered the next anchor in the instance started by the red anchor), but terminates as soon as the initial anchor is resolved.
@@
Figure 4. An illustration with four anchors in round $1$. Anchors follow the pre-assigned order: blue, red, green, pink. The blue, green and pink anchor commit directly; each has $\geq f+1$ references from nodes in round $2$. The red anchor, however, lacks support, and is only indirectly committed upon directly committing the yellow anchor in round $4$. The green and pink anchors must wait to be ordered until the red anchor is resolved.
@@
While simple, a straightforward instantiation of this design is not robust, and may, in practice negate (or even hurt) any latency benefit. By treating every node as anchor, progress is, in particular, bottlenecked on the slowest anchor. We illustrate such an unwelcome scenario in Figure 4: although the green and pink anchors have sufficient support to commit, they must wait until the preceding red anchor becomes committed or skipped. Unfortunately, the red anchor is slow and does not have any immediate links; it is indirectly committed only two rounds later via the next anchor in it’s consensus instance (yellow anchor). In even worse circumstances, an anchor might never commit. In this case, we must wait for confirmation that the anchor is skipped, and fill its assigned slot with a no-op. Such scenarios are not fringe outliers, but expected by design: since DAG nodes include only $12f+$ 1 edges, rounds are privy to advancing based on the fastest $12f+$ 1 nodes, and leaving the remainder behind. Thus, with high probability, at least one anchor per round will not receive sufficient direct votes.
@@
Adopting the reputation mechanism proposed by Shoal can help ameliorate, but not remedy the concern. We leverage reputation and modify Get_anchors(r) to limit the set of eligible anchors per round only to those associated with historically fast and well-connected replicas. Unfortunately, we find that in practice leveraging reputation alone is insufficient, and does not reliably enough exclude slow replicas. While past performance is a helpful indicator, network instabilities in geo-distributed systems make it hard to accurately predict future performance, and we observe that the set of ”slow” replicas changes very dynamically. To improve robustness further, we add two additional techniques that, respectively, try to (i) avoid unnecessarily classifying nodes as slow, and (ii) dynamically skip likely obsolete anchors.
Round Timeouts.
@@
Skipping Anchor Candidates.
Although round timeouts increase the reliability of the Direct Commit Rule, they cannot guarantee its success. Occasionally, replicas may fail or experience high networking delays, resulting in missing, or unconnected nodes. In such cases, all subsequently pre-ordered anchors must wait until the current anchor is resolved. Recall that an anchor $A1$ in round $r$ is only resolved only upon committing, in its local consensus instance, an anchor $A2$ in a future round $r^{\prime}>r$: $A1$ is committed if it is present in the causal history of $A2$ (Indirect Commit), or skipped if it is not. In practice, we expect indirectly committed anchors to be rare. Anchors that are merely slow, are likely to be committed thanks to the round timeout. Crashes, in turn, predominantly result in skips. Byzantine abuse can unfortunately not be avoided, and we must rely on the reputation scheme to avoid such anchors; fortunately, such faults are exceedingly rare in most practical deployments. [^4] Popular blockchain systems (Team, 2023, 2022) report observing zero Byzantine faults in over a year of deployment. Reported faults are typically benign crashes.
@@
5.3. Operating multiple DAGs in parallel
@@ 5.4. Discussion
We note that Shoal++ uses a linear star-based communication pattern to ceritfy nodes, while protocols such as PBFT (Castro and Liskov, 2002) leverages quadratic all-to-all communication. If desired, Shoal++ can adopt all-to-all communication as well, reducing latency by 1md (to 3.5md total) at the cost of increased message complexity. Finally, we remark that although Shoal++ incurs a small additional queuing latency, every replica acts as proposer, and clients need only contact their local replica (in absence of faults). Single leader designs, in contrast, require clients to contact possibly remote leaders, introducing additional ”queuing” latency.
6. Correctness
We now prove the Safety and Liveness of Shoal++. We separately prove each augmentation Shoal++ introduces through a reduction to the Safety and Liveness proofs of Bullshark and Shoal. To formalize the arguments, our proof relies on the following three properties.
@@ 6. Correctness
This holds true for all Narwhal-based DAGs, and follows directly from the fact that all nodes in the DAG are certified.
We next summarize the Safety guarantee of Bullshark.
@@
6.1. Fast Direct Commit Rule
@@
6.2. Multiple Anchors per Round
@@
By the lemma assumption and Property 3, anchors.pop() returns the same anchor in the first iteration of the while loop to all replicas (Line 12). By Lemma 6.1, all correct replicas agree on which anchor to order (Line 13). It follows, from the determinism of next_ordered_nodes and P.1 that all replicas return the same vector of nodes to be ordered in their first next_ordered_nodes call (line 18). In addition, by the determinism of next_ordered_nodes, all replicas end up with the same state (round and anchors) after the calls return. Therefore, by applying the argument above inductively, it is easy to show that all replicas also return the same vector of nodes in all future calls.
6.3. Queuing latency
Theorem 6.3.
@@ 7. Practical Considerations
Narwahl (Danezis et al., 2022) proposes, in addition to the DAG construction, to further decouple, and horizontally scale data dissemination using an additional ”worker”-layer. Workers disseminate batches optimistically a-priori, and DAG proposals contain only hash references. This construction is extremely throughput-friendly in the best case, as data is streamed continuously, and consensus carries only metadata. In practice, however, much like with uncertified DAG-BFT protocols (Keidar et al., 2022; Babel et al., 2023), other replicas might not have the data behind the hashes (or the references in the case of uncertified DAGs) when the proposal is received and thus additional 2md latency is required for fetching. Shoal++ opts to forgo the worker layer in order to avoid this latency, and chooses to disseminate transaction batches inline with DAG proposals. Operating staggered DAGs allows Shoal++ to nonetheless achieve throughput competitive to that of Narwahl (with a single worker), as smaller batches are sent more frequently, thus amortizing bandwidth close to a streaming design. [^6] Scaling throughput by employing more workers is typically not worth it in existing BFT deployments, as single-worker ordering throughput in existing protocols already far exceeds execution bandwidth. Finally, we note that inline data streaming simplifies data fetching as it avoids an additional indirection needed to dereference digests.
@@
Distance-based priority broadcast.
@@ 8. Evaluation
We implemented Shoal++ in the Aptos Blockchain codebase. Our prototype is written in Rust and utilizes the Tokio asynchronous runtime. It uses BLS (Boneh et al., 2001) implemented over BLS12-381 curves for signatures, RocksDB for persistent storage of consensus data, and Noise for authentication. We disable the execution and ledger storage components of the original blockchain in order to isolate consensus performance.
Figure 5. No failures with 100 replicas. Each data point represents the 50 percentile with error bars representing the 25, and 75th percentiles.
Baselines.
We compare Shoal++ against three popular, high throughput DAG-BFT consensus protocols, Bullshark (Spiegelman et al., 2022b), Shoal (Spiegelman et al., 2023), and Mysticeti (Babel et al., 2023), as well as one traditional low-latency BFT protocol, Jolteon (Gelashvili et al., 2022). Bullshark and Shoal represent state-of-the-art certified DAG-based protocols, while Mysticeti (Babel et al., 2023) is a concurrent work to Shoal++ proposing an uncertified design to reduce latency. Jolteon (Gelashvili et al., 2022) is a state-of-the-art leader-based BFT protocol, which improves Hotstuff (Yin et al., 2019) latency by 50%. Bullshark and Jolteon, respectively, are currently deployed on the Sui (Team, 2023) and Aptos (Team, 2022) Blockchains.
For an apples-to-apples comparison, we re-implemented Bullshark, Shoal, and Jolteon according to the papers’ description in the same codebase as Shoal++. Like Shoal++, Bullshark and Shoal are run without the Narwhal worker layer to minimize latency. For Mysticeti we run the publicly available source code; its prototype too is written in Rust, using Tokio, but forgoes writing consensus data to persistent storage, making it less production-realistic.
@@
We use the Google Cloud Platform (gcp, [n. d.]) to mimic the deployment of a globally decentralized network. Our testbed consists of 100 replicas spread evenly across 10 regions around the world: we choose 2 regions in the US (us-west1 and us-east1), 2 regions in Europe (europe-west4 and europe-southwest1), 1 region each in South America (southamerica-east1), South Africa (africa-south1) and Australia (australia-southeast1), and 3 in Asia (asia-northeast3, asia-southeast1, and asia-south1). The round-trip times between these regions range between 25ms and 317ms. We use n2d-standard-64 virtual machines, containing 64 vCPUs and 256 GB of memory [^7] https://cloud.google.com/compute/docs/general-purpose-machines#n2d_machines. We provisioned a 2TB network attached disk to each machine to guarantee enough IOPS for persistence. This spec is similar to those used by production blockchains and qualifies as commodity grade.
@@
Each client connects to a single replica and issues a continuous stream of dummy transactions, each consisting of 310 random bytes. The e2e latency is measured as the time between when a transaction arrives at a replica, and when the transaction is ordered and ready to be executed. We use a maximum batch size of 500 transactions in all systems; in all DAG-based systems, each DAG node proposal contains one batch, and for Jolteon the leader may propose up to 100 of them. We configure Jolteon to use a 1.5s timeout for view changes (the standard in existing production systems (Team, 2022)). We use a timeout of 600ms in each round for Bullshark, Shoal and Shoal++; Mysticeti by default uses a 1s round timeout. We note that timeouts in Bullshark, Shoal and Shoal++ serve different functions. Bullshark requires them for liveness, while Shoal manages to sidestep using timeouts for liveness all but edge cases; Shoal++ follows Shoal, but re-introduces a timeout only for improved performance. We note that the timeout is started at the beginning of a round, and thus typically introduces only a small additional wait past observing the first $12f+$ 1 certified nodes.
8.1. Failure-free case

### v2

@@ 4 Shoal++ Overview
More anchors
Next, Shoal++ takes Shoal’s idea of increased anchor frequency one step further, and tries to turn as many nodes into anchors as possible. Doing so requires striking balance. On the one hand, increasing the number of anchors increases the opportunities to avoid incurring the anchoring latency. On the other hand, it introduces also opportunity for more uncommitted anchors, which can stall progress. Adopting the reputation scheme proposed by Shoal can help identify a reliable subset of candidates, but is not enough to streamline progress in practice. Shoal++ (i) slows down round advancement to put the DAG into lockstep and avoid missed anchors, and (ii) dynamically re-interprets the tentative anchor schedule to skip anchors that are no longer needed.
More DAGs
Finally, Shoal++ minimizes Queuing latency by operating not one, but multiple DAGs in parallel. DAGs are staggered using a small offset, allowing waiting transactions that miss a round to quickly be included in the next DAG. This simultaneously improves throughput via better resource utilization: proposals are sent more frequently, batches are smaller and can be sent and processed more efficiently (in a streaming fashion) than less frequent, large batches.
5 Shoal++ In-Depth
Shoal++ uses as starting points the DAG-core of Narwahl [18], and Consensus-core of Bullshark [34] outlined in §3. For the sake of brevity, we do not duplicate the pseudocode that is available in previous works and, in particular, omit a detailed description of the DAG construction (i.e., node broadcasts, transaction submission, node fetching, garbage collection, and so on). We defer to the original papers for an in-depth discussion and correctness proofs.
Algorithm 1 summarizes the DAG API and node structure. We abstract away the node certification process, and the commit logic. Process_node and Process_certified_node respectively involve checking that a node proposal is valid and voting for it, and adding a new confirmed node to the DAG, triggering commit rules. Run_bullshark performs an instance of the Bullshark commit logic to determine (given a starting point) the first anchor to be ordered (§3.1.1).
Upon committing an anchor, all of its (yet-to-be-ordered) causal history is deterministically ordered as well. Helper functions causal_history and get_anchors aid, respectively, in identifying the causal history of a node (necessary for both commit and ordering logic), and the anchor candidate schedule. Shoal++ adopts and extends Shoal’s [33] reputation logic; get_anchors returns, for each round, a vector of eligible anchors.
@@ 5 Shoal++ In-Depth
4:     causal_history(v: Node) $\rightarrow$ Vec<Node>
@@
6:     get_anchors(round: int) -> Vec<Node>
@@
3:     weak_votes: Vec< Vec<int> > $\triangleright$ Initially empty
@@
5:     anchors: Vec<Node> $\triangleright$ Initially empty
@@
5.1 A faster Direct Commit rule
@@
Figure 3: Bullsharks’s direct commit rule requires f+1 certified vote nodes, while Shoal++’s direct commit rule requires $12f+$ 1 (uncertified) node proposals.
@@
We note, however, that this rule places a stronger requirement for liveness; $f$ additional proposals are (tentatively) required to link to an anchor. We thus retain the existing Direct Commit rule as backup, and allow replicas to commit using whichever rule is satisfied first. In some cases, it may be faster (and more reliable) to commit using $f+1$ certified nodes. For instance, when the network is unstable or point latencies are asymmetric, the fastest $f+1$ replicas might advance much more quickly than the fastest $12f+$ 1.
@@
In our empirical evaluation, we find that this case is rare (§8) and that replicas are, most of the time, able take advantage of the Fast Direct Commit rule. Shoal’s reputation scheme ensures that eligible anchor candidates are typically well connected, making it exceedingly commonplace for future round proposals to link to an anchor.
5.2 Increasing anchor frequency
@@
In order to reduce Anchoring latency, Shoal++ tries to dynamically designate as many nodes as possible as anchors. Intuitively, if all nodes were anchors that become committed, no node would experience any Anchoring latency. [^3] We note that such an approach is briefly discussed in the closing remarks of Shoal [33], but is neither implemented nor carefully evaluated for practical considerations. Safely operating more than one anchor per round requires additional orchestration. To ensure a consistently ordered log across all correct replicas, parallel anchors must be committed in a pre-defined order. One easy way to visualize parallel orchestration is by mapping anchors to fixed slots; much like parallel proposals in traditional BFT protocols such as PBFT [14]. Figure 4 shows a simple example in which there are four anchors in one round (all $n$ nodes are anchors). Each anchor employs it’s own consensus instance that may proceed in parallel, and out of order, but anchors can be committed only in sequence. Each consensus instance operates a one-shot Bullshark instance, augmented with the Fast Direct Commit Rule. We remark that each Bullshark instance operates it’s own tentative anchor schedule (e.g. the yellow anchor in Fig. 4 is considered the next anchor in the instance started by the red anchor), but terminates as soon as the initial anchor is resolved.
@@
Figure 4: An illustration with four anchors in round $1$, using the pre-assigned order: blue, red, green, pink. The blue anchor commits directly; it has $\geq f+1$ references from nodes in round $2$. The red anchor is only indirectly committed upon directly committing the yellow anchor in round $4$. The green and pink anchors must wait to be ordered until the red anchor is resolved.
@@
While simple, a straightforward instantiation of this design is not robust, and may, in practice negate (or even hurt) any latency benefit. By treating every node as an anchor, progress is, in particular, bottlenecked on the slowest anchor. We illustrate such an unwelcome scenario in Figure 4: although the green and pink anchors have sufficient support to commit, they must wait until the preceding red anchor becomes committed or skipped. Unfortunately, the red anchor is slow and does not have any immediate links; it is indirectly committed only two rounds later via the next anchor in it’s consensus instance (yellow anchor). In worse circumstances, an anchor might never commit. In this case, we must wait for confirmation that the anchor is skipped, and fill its assigned slot with a no-op. Such scenarios are not fringe outliers, but expected by design: since DAG nodes include only $12f+$ 1 edges, rounds are privy to advancing based on the fastest $12f+$ 1 nodes, and leaving the remainder behind. Thus, with high probability, at least one anchor per round will not receive sufficient direct votes.
@@
Adopting the reputation mechanism proposed by Shoal can help ameliorate, but not remedy the concern. We leverage reputation and modify Get_anchors(r) to limit the set of eligible anchors per round only to those associated with historically fast and well-connected replicas. Unfortunately, we find that in practice leveraging reputation alone is insufficient, and does not reliably enough exclude slow replicas. While past performance is a helpful indicator, network instabilities in geo-distributed systems make it hard to accurately predict future performance, and we observe that the set of "slow" replicas changes very dynamically. To improve robustness further, we add two additional techniques that, respectively, try to (i) avoid unnecessarily classifying nodes as slow, and (ii) dynamically skip likely obsolete anchors.
Round Timeouts
@@
Skipping Anchor Candidates
Although round timeouts increase the reliability of the Direct Commit Rule, they cannot guarantee its success. Occasionally, replicas may fail or experience high networking delays, resulting in missing, or unconnected nodes. In such cases, all subsequently pre-ordered anchors must wait until the current anchor is resolved. Recall that an anchor $A1$ in round $r$ is only resolved only upon committing, in its local consensus instance, an anchor $A2$ in a future round $r^{\prime}>r$: $A1$ is committed if it is present in the causal history of $A2$ (Indirect Commit), or skipped if it is not. In practice, we expect indirectly committed anchors to be rare. Anchors that are merely slow, are likely to be committed thanks to the round timeout. Crashes, in turn, predominantly result in skips. Byzantine abuse can unfortunately not be avoided, and we must rely on the reputation scheme to avoid such anchors; fortunately, such faults are exceedingly rare in most practical deployments. [^4] Blockchain systems [43, 41] report observing zero Byzantine faults in over a year of deployment. Reported faults are typically benign crashes.
@@
5.3 Operating multiple DAGs in parallel
@@ 5.4 Discussion
We note that Shoal++ uses a linear star-based communication pattern to ceritfy nodes, while protocols such as PBFT [14] leverages quadratic all-to-all communication. If desired, Shoal++ can adopt all-to-all communication as well, reducing latency by 1md (to 3.5md total) at the cost of increased message complexity. Finally, we remark that although Shoal++ incurs a small additional queuing latency, every replica acts as proposer, and clients need only contact their local replica (in absence of faults). Single leader designs, in contrast, require clients to contact possibly remote leaders, introducing additional "queuing" latency.
6 Correctness
We now prove the Safety and Liveness of Shoal++. We separately prove each augmentation Shoal++ introduces through a reduction to the Safety and Liveness proofs of Bullshark [34] and Shoal [33]. To formalize the arguments, our proof relies on the following three properties:
@@ 6 Correctness
Property 1 holds true for all Narwhal-based DAGs, and follows directly from the fact that all nodes in the DAG are certified; Properties 2 and 3 follow, respectively, from the safety proofs of Bullshark and Shoal.
Fast Direct Commit Rule
@@
Multiple Anchors per Round
@@
By the lemma assumption and Property 3, anchors.pop() returns the same anchor in the first iteration of the while loop to all replicas (Line 12). By Lemma 1, all correct replicas agree on which anchor to order (Line 13). It follows, from the determinism of next_ordered_nodes and P.1 that all replicas return the same vector of nodes to be ordered in their first next_ordered_nodes call (line 18). In addition, by the determinism of next_ordered_nodes, all replicas end up with the same state (round and anchors) after the calls return. Therefore, by applying the argument above inductively, it is easy to show that all replicas also return the same vector of nodes in all future calls. ∎
Theorem 1.
@@ 7 Practical Considerations
Narwahl [18] proposes, in addition to the DAG construction, to further decouple, and horizontally scale data dissemination using an additional "worker"-layer. Workers disseminate batches optimistically a-priori, and DAG proposals contain only hash references. This construction is extremely throughput-friendly in the best case, as data is streamed continuously, and consensus carries only metadata. In practice, however, much like with uncertified DAG-BFT protocols [26, 11], other replicas might not have the data behind the hashes (or the references in the case of uncertified DAGs) when the proposal is received and thus additional 2md latency is required for fetching. Shoal++ opts to forgo the worker layer in order to avoid this latency, and chooses to disseminate transaction batches inline with DAG proposals. Operating staggered DAGs allows Shoal++ to nonetheless achieve throughput competitive to that of Narwahl (with a single worker), as smaller batches are sent more frequently, thus amortizing bandwidth close to a streaming design. [^6] Scaling throughput by employing more workers is typically not worth it in existing BFT deployments, as single-worker ordering throughput in existing protocols already far exceeds execution bandwidth. Finally, we note that inline data streaming simplifies data fetching as it avoids an additional indirection needed to dereference digests.
@@
Distance-based priority broadcast
@@ 8 Evaluation
We implemented Shoal++ into an open-source code base of Aptos, one of the major blockchain projects. Our prototype is written in Rust and utilizes the Tokio [9] asynchronous runtime. It uses BLS [13] implemented over BLS12-381 curves for signatures, RocksDB [8] for persistent storage of consensus data, and Noise [7] for authentication. We disable the execution and ledger storage components of the original blockchain in order to isolate consensus performance. Each data point represents the 50 percentile (median) with error bars representing the 25, and 75th percentiles.
Figure 5: No failures with 100 replicas.
Baselines
We compare Shoal++ against three popular high throughput DAG-BFT consensus protocols, Bullshark [35], Shoal [33], and Mysticeti [11], as well as one traditional low-latency BFT protocol, Jolteon [21]. Bullshark and Shoal represent state-of-the-art certified DAG-based protocols, while Mysticeti [11] is a concurrent work to Shoal++ that proposes an uncertified design to reduce latency. Jolteon [21] is a state-of-the-art leader-based BFT protocol, which improves Hotstuff [46] latency by 50%. A variant of Jolteon is currently deployed on Aptos [41]. Sui [43] recently replaced Bullshark with a Mysticeti deployment. For an apples-to-apples comparison, we re-implemented Bullshark, Shoal, and Jolteon according to the papers’ description in the same codebase as Shoal++. For Mysticeti we run the publicly available source code [6] referred to in the latest version of the paper; its prototype too is written in Rust, using Tokio, but forgoes writing consensus data to persistent storage, making it less production-realistic. To minimize latency, we opt to disable the Narwhal [18] worker layer for both Bullshark and Shoal. However, for a fair throughput comparison with Shoal++, we additionally augment both systems with Shoal++’s proposed parallel-DAG strategy ("Bullshark/Shoal More DAGs", Fig. 5).
@@
We use the Google Cloud Platform [4] to mimic the deployment of a globally decentralized network. Our testbed consists of 100 replicas spread evenly across 10 regions around the world: we choose two regions in the US (us-west1 and us-east1), two in Europe (europe-west4 and europe-southwest1), three in Asia (asia-northeast3, asia-southeast1, and asia-south1), and, respectively, one each in South America (southamerica-east1), South Africa (africa-south1) and Australia (australia-southeast1). The round-trip times between these regions range between 25ms and 317ms. We use n2d-standard-64 virtual machines, containing 64 vCPUs and 256 GB of memory [3]. We provisioned a 2TB network attached disk to each machine to guarantee enough IOPS for persistence. This spec is similar to those used by production blockchains and qualifies as commodity grade.
@@
Clients connect to a single (local) replica and issue a continuous stream of dummy transactions (310 random bytes). As is standard in blockchain deployments, we measure consensus latency as the time between when a transaction first arrives at a replica, and when the transaction is ordered and ready to be executed. We use a batch size of 500 transactions across all systems. In all DAG-based systems, each node proposal contains one batch; Jolteon proposals may contain up to 100 batches. We configure Jolteon to use a 1.5s timeout for view changes (the standard in existing production systems [41]), and use a round timeout of 600ms for Bullshark, Shoal and Shoal++; Mysticeti by default uses a 1s round timeout. We remark that timeouts in Bullshark, Shoal and Shoal++ serve different functions. Bullshark requires them for liveness, while Shoal manages to sidestep using timeouts for liveness all but edge cases; Shoal++ follows Shoal, but re-introduces a timeout only for improved performance. We note that the timeout is started at the beginning of a round, and thus typically introduces only a small additional wait past observing the first $12f+$ 1 certified nodes.
8.1 Failure-free case